Muller Corporation purchased a new truck on January 1, XXX1 for $55,000 cash. Muller estimated the truck would have a salvage value of $4,000 at the end of the useful life of 5 years (i.e. at the end of Year XXX5). On January 1, XXX3, Muller had to replace the engine of the truck paying $7,500 cash. Due to the replacement of the engine, Muller estimates that the truck will continue to be productive for another four years (i.e. until the end of year XXX6) and the estimated salvage value still remains the same at the end of the productive life. The company decides to capitalize the cost of replacing the engine since the productive life is extended by a year because of the new engine. Required: (a) Assuming straight-line depreciation is used, Calculate the depreciation expense for year XXX1 and for Year XXX5. (b) Suppose, Muller sells the truck on January 1, XXX6 for $15,000, estimate the impact of that sale on the after-tax income reported in the income statement. Assume an income tax rate of 30%.
Answer #1

a. Depreciation expense for year XXX1 = $ ( 55,000 - 4,000) / 5 = $ 10,200.

Depreciation expense for year XXX5 = $ [ ( 55,000 - 10,200 x 2 ) - 4,000] / 4 ] + $ 7,500 / 4 = $ 9,525

b. Book value of the truck on January 1, XXX6 = $ ( 55,000 + 7,500) - ($ 10,200 x 2) - ( 9,525 x 3 ) = $ 13,525

Gain on sale of truck = $ 15,000 - $ 13,525 = $ 1,475.

After tax income will increase by ( 9,525 - 1,475) x 0.70 = $ 5,635